OpenStudy (anonymous):

You plan a car trip for which you want to average 90 km/h. You cover the first half of the distance at an average speed of only 48 km/h. What must your average speed be in the second half of the trip to meet your goal? Not that the velocities are based on half the distance, not half the time.

OpenStudy (anonymous):

720 km/h Better hurry...

OpenStudy (anonymous):

how did you figure this out?

OpenStudy (anonymous):

Say S is half the distance, V1, V2 and Vtot the speeds, T1, T2 and Ttot the times.

OpenStudy (anonymous):

You know V1=48 km/h and Vtot=90km/h

OpenStudy (anonymous):

T1=S/V1 T2=Ttot-T1=(2S/Vtot)-S/V1 V2=S/T2 = S / (2S/Vtot-S/V1)

OpenStudy (anonymous):

Simplify in V2 = V1Vtot/(2V1-Vtot) = 48 * 90 / 6 = 720

OpenStudy (anonymous):

Interestingly, there is only a real solution if V1>45. For V1=45, you're only half way in the time you need to get to you destination (no matter how far it is)